OpenStudy (anonymous):

Find x and y such that : lnx.lny=6 and xy=e^5.

OpenStudy (anonymous):

from the first equation lnx =6/ln y------------(1) taking ln for the second equation ln x +ln y =5 substituting (1) in the above equation you get , (6/ln y) + ln y = 5 solve for ln y and substitute the ans into the first equation and get the value of ln x and then take anti log

OpenStudy (anonymous):

(lnx)(lny) = 6 xy = e^5 This means ln(xy) = ln(e^5) ln(xy) = 5 lnx + lny = 5. (one of the laws of logarithms) Let w = lnx and v = lny, then the 2 equations become: wv = 6 w + v = 5. Those are easy to solve for w and v. Once you have w and v you can find x and y using w = lnx means e^w = x v = lny means e^v = y

OpenStudy (anonymous):

got it!!!

OpenStudy (anonymous):

Ok ! Good solution.

OpenStudy (anonymous):

I type fast. I didn't mean to scoop DHASHNI.

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

You were answering when I started to answer and I finished first, which was sort of rude and I didn't mean to be rude. That's all.

OpenStudy (anonymous):

k!!!